OpenStudy (anonymous):

Math Help Please!! Write the equation of the direct variation that includes the point (8, 18).

jimthompson5910 (jim_thompson5910):

"direct variation" means we'll have y = k*x for some fixed number k

jimthompson5910 (jim_thompson5910):

y = k*x 18 = k*8 k = ??

OpenStudy (anonymous):

Would K represent 9 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

k = 9 is false

OpenStudy (anonymous):

I'm not sure what K could be @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you divide both sides by 8 to isolate k

OpenStudy (anonymous):

18 divided by 8 is 2.25.

jimthompson5910 (jim_thompson5910):

leave it as a fraction

jimthompson5910 (jim_thompson5910):

reduce as much as possible

OpenStudy (anonymous):

What would be the fraction that I have to reduce?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

reduce 18/8 to get ??

OpenStudy (anonymous):

Y = 9/4

jimthompson5910 (jim_thompson5910):

yep, your equation is \[\Large y = \frac{9}{4}x\]

OpenStudy (anonymous):

Thank you so much! (:

jimthompson5910 (jim_thompson5910):

well you had x on there, but deleted it for some reason

jimthompson5910 (jim_thompson5910):

you're welcome